# Confidential Serverless Computing

### v1

@@ Abstract
We present Hacher, a confidential computing system for secure serverless deployments to overcome these limitations. By employing nested confidential execution and a decoupled guest OS within CVMs, Hacher runs each function in a minimal "trustlet", significantly improving security through a reduced Trusted Computing Base (TCB). Furthermore, by leveraging a data-centric I/O architecture built upon a lightweight LibOS, Hacher optimizes network communication to address performance and resource efficiency challenges.
Our evaluation shows that compared to CVM-based deployments, Hacher has 4.3 smaller TCB, improves end-to-end latency (15–93%), achieves higher function density (up to 907), and reduces inter-function communication (up to 27) and function chaining latency (16.7-30.2); thus, Hacher offers a practical system for confidential serverless computing.
@@ 1 Introduction
Serverless computing is a rapidly evolving cloud paradigm offering low costs and simplified deployment [68]. As serverless applications increasingly manage sensitive data, particularly in AI/ML workloads, strong security measures are of utmost importance [21, 19]. However, securely managing data within serverless computing is particularly challenging due to their ephemeral and distributed execution, where short-lived functions are chained through network communication [8, 147, 77].
Confidential Virtual Machines (CVMs) [75, 16, 17] provide a promising approach for the secure execution of serverless workloads in untrusted cloud environments [32], as they protect the confidentiality and integrity of entire VMs and do not require significant application modifications [74, 18].
@@
Despite the advancements in virtualization architectures, no existing approach can fully resolve the combined challenges of security, performance, and resource efficiency in confidential serverless computing. Traditional VMs provide hardware-based isolation but incur substantial boot delays [101], high inter-VM communication costs [151], and maintain large TCBs because of their full virtualization stacks. Lightweight VMs (e.g., FireCracker [3], Kata [41]) optimize function density and startup latency [100], and reduce TCBs by trimming hypervisor and guest OS components [3]; however, they still suffer from significant I/O overheads and designs that complicate security verification [40, 137, 143]. Specialized OS architectures (e.g., LibOSes [133, 26], Unikernels [85, 98]) deliver near-instant startups by compiling applications with minimal kernel libraries but require developer efforts to craft application-specific images and sacrifice compatibility. CVMs (e.g., AMD SEV-SNP [9], Intel TDX [75], ARM CCA [17]), in turn, provide hardware-based security properties but introduce significant network communication overheads [90, 127, 109]. In summary, while each approach addresses individual aspects of the problem, a holistic solution that reconciles all these concerns remains an open challenge.
@@
As a solution, we propose Hacher, a lightweight confidential computing system that enables secure serverless deployments in untrusted clouds by fundamentally rethinking how sensitive workloads are isolated and executed inside CVMs. In Hacher, each function runs in its own trustlet—a minimal serverless process that executes within a secure environment inside a CVM. With this design, client requests are securely routed through the untrusted guest OS, while the actual execution occurs in trustlets that encapsulate only the critical application code and the minimal system support required to serve these functions. This arrangement leverages two key mechanisms: nested confidential execution and a decoupled guest OS architecture. Inspired by nested virtualization [23] and ARM’s TrustZone model [18, 130], we partition the CVM space into smaller nested isolated regions. By confining sensitive code to trustlets, Hacher significantly reduces the overall TCB and minimizes the attack surface, thereby overcoming the limitations of hosting workloads on a monolithic guest OS.
@@
To achieve high performance and fast function instantiation, Hacher adopts several techniques. First, it employs pre-initialized process templates (inspired by Android’s zygote model [36, 87]) to expedite the bootstrapping of trustlets. Complementing this, a dynamically loadable LibOS enables trustlets to achieve the minimal startup latency characteristic of unikernel-based systems [98, 121, 161] while retaining flexibility for diverse workloads. Moreover, Hacher introduces a data-centric I/O architecture that leverages high-density function co-location and replaces conventional CVM networking for inter-function communication, optimizing communication paths via secure data objects. It also incorporates differential attestation-an innovative technique that reuses pre-measured components to reduce attestation latency by focusing on only the mutable parts during invocation.
@@
Collectively, these mechanisms enhance resource efficiency. As process templates and the LibOS are designed to share memory, and trustlets are ephemeral—occupying resources only during function execution—Hacher achieves a higher density of function deployments per CVM. This tight integration of performance optimizations with a thin TCB not only accelerates startup and execution but also allows for more efficient resource utilization, thereby supporting a larger number of concurrent function requests within the same memory footprint.
@@
We implement Hacher on AMD SEV-SNP [9] with source code and evaluation setup to be publicly available. Our evaluation shows Hacher reduces TCB by 4.3 compared to CVMs while achieving lower end-to-end latency in real serverless applications for both cold (85.10-93.44%) and warm (14.97%) starts. In large-scale simulations using Azure Functions traces [160, 139], Hacher achieves significantly lower invocation latency than CVMs (i.e., 5ms vs. 489s at the 50th percentile, 1.5s vs. 881s at the 99th percentile). Further, performance microbenchmarks indicate that Hacher’s cold starts are faster than CVMs and 35% than traditional VMs, while warm starts show negligible latency (<10.3 ms). Function communication latency decreases by thanks to Hacher’s data-centric networking architecture, while function chaining performance improves by over CVMs. Attestation latency is also reduced as Hacher minimizes the measurement time (0.25 ms in warm starts). Lastly, Hacher has a smaller memory footprint than CVMs (up to ) for the same number of concurrent functions, rendering Hacher an ideal solution for secure serverless deployments in untrusted clouds.
@@ 2 Confidential Virtual Machines (CVMs)
| Baseline | Host Kernel | Firm- ware | Guest Kernel | Runtime | Hyper- visor | Total |
| LibOS (Gramine) | 1,115 | 903 | — | 36 | — | 2,054 |
| Containers (Kata) | 1,115 | 903 | 1,809 | 8,001 | 1,757 | 13,585 |
| VM (KVM-Linux) | 1,115 | 903 | 3,177 | 744 | 1,757 | 7,696 |
| CVM (SEV-SNP) | — | 903 | 3,177 | 744 | — | 4,824 |
| Hacher | — | 332 | — | 780 | — | 1,112 |
@@ 3.1 Issue #1: Large Trusted Computing Base
Table 1 showcases that the CVM TCB size (4.8M LoC) is smaller than traditional VMs (7.7M LoC) but much larger than LibOS alternatives (e.g., Gramine: 2M LoC). Thus, CVM-hosted serverless stacks do not adequately minimize the TCB. The inclusion of a full guest OS increases complexity and exposes a large attack surface. In contrast, Hacher achieves approximately 4 TCB reduction compared to CVMs and 2 compared to Gramine without compromising security.
@@ 3.2 Issue #2: Slow Boot Times
To this end, we measure the boot time of different virtualization architectures (Figure 1), which indicate that CVMs exhibit a boot time of 8.3 seconds, 2.2 slower than standard VMs (3.7s), primarily due to VMM setup, OS initialization, and encrypted memory configuration. These inevitable security-induced boot delays often exceed function lifetimes, making CVMs impractical for latency-sensitive serverless computing. In contrast, with an initialized Hacher, we can achieve faster cold starts (2.4 s) than traditional VMs (3.7 s), and a near-negligible warm start latency (<10.3 ms).
@@ 3.3 Issue #3: High Networking Overheads
We further examine the network performance of CVMs with various Linux configurations. Figure 1 shows iPerf [76] throughput (UDP, 1460 bytes payload, client runs on the host). “swotlb” means that bounce buffer is enabled, and “vhost” implies using the vhost optimization. The SNP VM always uses a bounce buffer. We observe that the SNP VM suffers from performance degradation (only 17% of the VM), and the guest-side halt polling [45] (“-hpoll”) mitigates the performance overhead at the cost of excessive CPU usage [109]. In contrast, Hacher achieves lower latency than CVMs (1.4-27) leveraging function co-location and following a data-centric networking model for secure, yet performant, communication.
@@ 3.4 Issue #4: Inefficient Scheduling
Specifically, confidential serverless functions in isolated environments cannot share resources, restricting the flexibility of resource-aware schedulers [140]. The limited pool of instances per CVM node forces balancing instance reuse, increasing cold start ratios [163, 96]. In addition, switching execution context between CVMs incurs significant overheads (e.g., TLB/cache flushing, memory validation) [27, 109, 164]. Our large-scale system simulation using sampled [149] Azure Functions traces [160, 139] (30 minutes, 4,000 functions, 3.8 million invocations) shows CVMs having higher scheduling latency across all percentiles compared to VMs and microVMs (Figure 2) under the same scheduling model.
Thus, we conclude that standard scheduling is suboptimal for CVM-based serverless. Hacher achieves lower scheduling latency with better tail latencies by sustaining more warm instances per node and using a run-to-completion model [163].
@@ 3.5 Issue #5: Impractical Consolidation
Thus, CVM-based serverless struggles to consolidate functions efficiently. In contrast, Hacher can host multiple isolated functions within a CVM (up to 907 less memory than CVMs), exceeding the hardware-imposed density limit.
@@ 3.6 Issue #6: High Attestation Overheads
Figure 2 shows that the measurement of a full guest OS kernel for a Linux CVM can take up to 1.1 s, significantly increasing the overall startup latency. While long-running workloads can amortize this cost, the short execution times of serverless functions make the repetitive attestation overhead prohibitive. Our Hacher prototype significantly reduces the measurement cost to 0.25 ms for warm starts by reusing already calculated measurements across the same functions.
@@ 4 Overview
As a solution, we present Hacher, a lightweight confidential computing system for secure serverless deployments in untrusted cloud environments.
@@ 4.1 System Architecture
Hacher builds on CVMs and uses hardware-enforced privilege partitioning (e.g., VMPL [10], TD-partitioning [70]) to create lightweight trustlets — minimal serverless processes that execute within a secure environment inside a CVM. Figure 3 presents Hacher’s key components: the trusted monitor, the trusted processes, the data objects and the zygote registry.
The trusted monitor is a small, privileged, and nested software component within Hacher’s TCB. It manages the instantiation and invocation of serverless functions within lightweight protection domains isolated from the guest OS. It performs security-critical tasks such as memory management, function deployment, and attestation.
Within the CVM, Hacher introduces the notion of trusted processes for hosting serverless functions. These processes possess enclave-like features: (i) isolation from the guest OS, (ii) trusted boot, and (iii) remote attestation capabilities. They are categorized into zygotes and trustlets. Zygotes serve as pre-initialized templates containing runtime environments (e.g., Python), while trustlets are lightweight instances, derived from zygotes, to execute functions in an isolated manner. To facilitate inter-function communication, Hacher provides the abstraction of data objects, managed by the trusted monitor ensuring isolation. This design favors reduced boot times, efficient memory use via copy-on-write [5, 52, 113, 51, 94], and optimized data exchange between co-located functions.
@@ 4.2 Life of a Request
4.3 Threat Model
Hacher extends the CVM threat model [9, 29, 6, 124] by excluding the guest OS from its trust boundaries. We consider adversaries controlling the guest OS attempting to compromise serverless functions within trusted processes, the host OS attacking the CVM via VM-VMM interfaces (e.g., network I/O), or deploying malicious functions to access neighboring functions and leak data. Users trust function providers, but serverless and platform providers are potentially malicious.
The trusted monitor is Hacher’s only trusted software, enabling remote attestation. We assume that the platform hardware/firmware functions correctly and deployed functions do not intentionally leak data. Users must employ encrypted protocols (e.g., TLS) for communication. Hacher does not address denial-of-service, physical, and side-channel attacks.
@@ 4.4 Design Principles and Primitives
We now present the design principles and primitives of Hacher that address the six key problems, outlined in § 3.
#1: Nested confidential execution. As a countermeasure to the large TCB, the hardware-limited scalability and the costly inter-function communication of CVM-based serverless deployments, Hacher introduces nested confidential execution. Inspired by nested virtualization [23, 81, 31, 107, 159]—not directly applicable to CVMs due to host VMM support requirements [64, 4]—Hacher leverages intra-CVM isolation [1, 70, 108] to partition runtime components [102]. This allows the secure coexistence of untrusted components in lower-privilege compartments, while critical services operate in highly privileged domains, reducing the software TCB.

### v3

@@ Abstract
We present Wallet, a lightweight confidential computing system for secure serverless deployments. By employing nested confidential execution and a decoupled guest OS within CVMs, Wallet runs each function in a minimal “trustlet”, significantly improving security through a reduced Trusted Computing Base (TCB). Furthermore, by leveraging a data-centric I/O architecture built upon a lightweight LibOS, Wallet optimizes network communication to address performance and resource efficiency challenges.
Our evaluation shows that compared to CVM-based deployments, Wallet has a 4.3 smaller TCB, improves end-to-end latency (15–93%), achieves higher function density (up to 907), and reduces inter-function communication (up to 27) and function chaining latency (16.7-30.2); thus, Wallet offers a practical system design for confidential serverless computing.
@@ 1 Introduction
Serverless computing is a rapidly evolving cloud paradigm offering low costs and simplified deployment [70]. As serverless applications increasingly manage sensitive data, particularly in AI/ML, strong security measures are of utmost importance [22, 19]. However, securely managing data within serverless computing is particularly challenging due to their ephemeral and distributed execution, where short-lived functions are chained through network communication [8, 152, 79].
To this end, Confidential Virtual Machines (CVMs) [77, 16, 17] provide a promising approach for the secure execution of serverless workloads in untrusted cloud environments [33], as they protect the confidentiality and integrity of entire VMs and do not require significant application modifications [76, 18].
@@
Despite the advancements in virtualization architectures, no existing approach can fully resolve the combined challenges of security, performance, and resource efficiency in confidential serverless computing. Traditional VMs provide hardware-based isolation but incur substantial boot delays [103], high inter-VM communication costs [156], and maintain large TCBs because of their full virtualization stacks. Lightweight VMs (e.g., FireCracker [3], Kata [42]) optimize function density and startup latency [102], and reduce TCBs by trimming hypervisor and guest OS components [3]; however, they still suffer from significant I/O overheads and designs that complicate security verification [41, 140, 148]. Specialized OS architectures (e.g., LibOSes [136, 27, 88], Unikernels [87, 100, 111]) deliver near-instant startups by compiling applications with minimal kernel libraries but require developer efforts to craft application-specific images and sacrifice compatibility. CVMs (e.g., AMD SEV-SNP [9], Intel TDX [77], ARM CCA [17]), in turn, provide hardware-based security properties but introduce significant network communication overheads [92, 130, 112]. In summary, while each approach addresses individual aspects of the problem, a holistic solution that reconciles all these concerns remains an open challenge.
@@
As a solution, we propose Wallet, a lightweight confidential computing system that enables secure serverless deployments in untrusted clouds by fundamentally rethinking how sensitive workloads are isolated and executed inside CVMs. In Wallet, each function runs in its own trustlet—a minimal serverless process that executes within a secure environment inside a CVM. With this design, client requests are securely routed through the untrusted guest OS, while the actual execution occurs in trustlets that encapsulate only the critical application code and the minimal system support required to serve these functions. This arrangement leverages two key mechanisms: nested confidential execution and a decoupled guest OS architecture. Inspired by nested virtualization [24] and ARM’s TrustZone model [18, 133], we partition the CVM space into smaller nested isolated regions. By confining sensitive code to trustlets, Wallet significantly reduces the overall TCB and minimizes the attack surface, thereby overcoming the limitations of hosting workloads on a monolithic guest OS.
@@
To achieve high performance and fast function instantiation, Wallet adopts several techniques. First, it employs pre-initialized process templates (inspired by Android’s zygote model [37, 89]) to expedite the bootstrapping of trustlets. Complementing this, a dynamically loadable LibOS enables trustlets to achieve the minimal startup latency characteristic of unikernel-based systems [100, 124, 166] while retaining flexibility for diverse workloads. Moreover, Wallet introduces a data-centric I/O architecture that leverages high-density function co-location and replaces conventional CVM networking for inter-function communication, optimizing communication paths via secure data objects. It also incorporates differential attestation-an innovative technique that reuses pre-measured components to reduce attestation latency by focusing on only the mutable parts during invocation.
@@
Collectively, these mechanisms enhance resource efficiency. As process templates and the LibOS are designed to share memory, and trustlets are ephemeral—occupying resources only during function execution—Wallet achieves a higher density of function deployments per CVM. This tight integration of performance optimizations with a thin TCB not only accelerates startup and execution but also allows for more efficient resource utilization, thereby supporting a larger number of concurrent function requests within the same memory footprint.
@@
We implement Wallet on AMD SEV-SNP [9] with source code and evaluation setup to be publicly available. Our evaluation shows Wallet reduces TCB by 4.3 compared to CVMs while achieving lower end-to-end latency in real serverless applications for both cold (85.10-93.44%) and warm (14.97%) starts. In large-scale simulations using Azure Functions traces [165, 142], Wallet achieves significantly lower invocation latency than CVMs (i.e., 5ms vs. 489s at the 50th percentile, 1.5s vs. 881s at the 99th percentile). Further, performance microbenchmarks indicate that Wallet’s cold starts are faster than CVMs and 35% than traditional VMs, while warm starts show negligible latency (<10.3 ms). Function communication latency decreases by thanks to Wallet’s data-centric networking architecture, while function chaining performance improves by over CVMs. Attestation latency is also reduced as Wallet minimizes the measurement time (0.25 ms in warm starts). Lastly, Wallet has a smaller memory footprint than CVMs (up to ) for the same number of concurrent functions, rendering Wallet an ideal solution for secure serverless deployments in untrusted clouds.
@@ 2 Confidential Virtual Machines (CVMs)
Currently, Linux supports Intel TDX, AMD SEV-SNP, and ARM CCA (the hardware support of ARM CCA is absent as of now). In addition, Gramine-TDX [88] minimizes the guest TCB by employing a LibOS architecture and specializing the OS to a specific application on the Intel TDX platform.
@@ 3.1 Issue #1: Large Trusted Computing Base
Table 1 showcases that the CVM TCB size (4.8M LoC) is smaller than traditional VMs (7.7M LoC) but much larger than LibOS alternatives (e.g., Gramine LibOS: 2M LoC). Thus, CVM-hosted serverless stacks do not adequately minimize the TCB. The inclusion of a full guest OS increases complexity and exposes a large attack surface. In contrast, Wallet achieves approximately 4 TCB reduction compared to CVMs and 2 compared to Gramine LibOS without compromising security.
@@ 3.2 Issue #2: Slow Boot Times
To this end, we measure the boot time of different virtualization architectures (Figure 1), which indicate that CVMs exhibit a boot time of 8.3 seconds, 2.2 slower than standard VMs (3.7s), primarily due to VMM setup, OS initialization, and encrypted memory configuration. These inevitable security-induced boot delays often exceed function lifetimes, making CVMs impractical for latency-sensitive serverless computing. While Wallet has overall the highest initialization time ( 12s), it is an one-time cost. In contrast, with an initialized Wallet, we can achieve faster cold starts (2.4 s) than traditional VMs (3.7 s), and a near-negligible warm start latency (<10.3 ms). Notably, this warm start achieves even lower start-up times compared to the LibOS ( 28 ms) on its own running without isolation on the host, thanks to its efficient CoW-based function execution.
@@ 3.3 Issue #3: High Networking Overheads
We further examine the network performance of CVMs with various Linux configurations. Figure 1 shows iPerf [78] throughput (UDP, 1460 bytes payload, client runs on the host). “swotlb” means that bounce buffer is enabled, and “vhost” implies using the vhost optimization. The SNP VM always uses a bounce buffer. We observe that the SNP VM suffers from performance degradation (only 17% of the VM), and the guest-side halt polling [46] (“-hpoll”) mitigates the performance overhead at the cost of excessive CPU usage [112]. In contrast, Wallet achieves lower latency than CVMs (1.4-27) leveraging function co-location and following a data-centric networking model for secure, yet performant, communication.
@@ 3.4 Issue #4: Inefficient Scheduling
Specifically, confidential serverless functions in isolated environments cannot share resources, restricting the flexibility of resource-aware schedulers [144]. The limited pool of instances per CVM node forces balancing instance reuse, increasing cold start ratios [168, 98]. In addition, switching execution context between CVMs incurs significant overheads (e.g., TLB/cache flushing, memory validation) [28, 112, 169]. Our large-scale system simulation using sampled [154] Azure Functions traces [165, 142] (30 minutes, 4,000 functions, 4.1 million invocations) shows CVMs having higher scheduling latency across all percentiles compared to VMs and microVMs (Figure 2) under the same scheduling model.
Thus, we conclude that standard scheduling is suboptimal for CVM-based serverless. Wallet achieves lower scheduling latency with better tail latencies by sustaining more warm instances per node and using a run-to-completion model [168].
@@ 3.5 Issue #5: Impractical Consolidation
Thus, CVM-based serverless struggles to consolidate functions efficiently. In contrast, Wallet can host multiple isolated functions within a CVM (up to 907 less memory than CVMs), exceeding the hardware-imposed density limit.
@@ 3.6 Issue #6: High Attestation Overheads
Figure 2 shows that the measurement of a full guest OS kernel for a Linux CVM can take up to 1.1 s, significantly increasing the overall startup latency. While long-running workloads can amortize this cost, the short execution times of serverless functions make the repetitive attestation overhead prohibitive. Our Wallet prototype significantly reduces the measurement cost to 0.25 ms for warm starts by reusing already calculated measurements across the same functions.
@@ 3.7 Summary and Problem Statement
To this end, there exists prior work that focuses on providing specialized solutions to tackle these challenges [88, 65]. Specifically, Gramine-TDX [88] provides a slim, security-first OS kernel, based on Gramine LibOS [153], to run unmodified applications in CVMs while having a minimal TCB. However, it is mainly suited for single-tenant architectures, where each serverless function requires its dedicated CVM, which conflicts with confidential serverless computing requirements.
Precisely, Gramine-TDX cannot achieve efficient function consolidation due to hardware CVM limits, requires per-function attestation, and forces inter-function communication through expensive CVM networking. Further, Gramine-TDX cannot leverage function co-location opportunities within the same CVM in a secure manner, as it operates with system-wide privileges and lacks the fine-grained privilege separation mechanisms necessary to safely isolate multiple co-located functions from each other and from the underlying software layers. These architectural constraints make Gramine-TDX unsuitable for the multi-tenant, high-density, and communication-intensive nature of serverless workloads.
@@ 4 Overview
As a solution, we present Wallet, a lightweight confidential computing system for secure serverless deployments in untrusted cloud environments.
4.1 Threat Model
Wallet extends the CVM threat model [9, 30, 6, 127] by excluding the guest OS from its trust boundaries. We consider adversaries controlling the guest OS attempting to compromise serverless functions within trusted processes, the host OS attacking the CVM via VM-VMM interfaces (e.g., network I/O), or deploying malicious functions to access neighboring functions and leak data. Users trust function providers, but serverless and platform providers are potentially malicious.
The trusted monitor is Wallet’s only trusted software, enabling remote attestation. We assume that the platform hardware/firmware functions correctly and deployed functions do not intentionally leak data. Users must employ encrypted protocols (e.g., TLS) for communication. Wallet does not address denial-of-service, physical, and side-channel attacks.
@@ 4.2 System Architecture
Wallet builds on CVMs and uses hardware-enforced privilege partitioning (e.g., VMPL [10], TD-partitioning [72]) to create lightweight trustlets — minimal serverless processes that execute within a secure environment inside a CVM. Figure 3 presents Wallet’s key components: the trusted monitor, the trusted processes, the data objects and the zygote registry.
The trusted monitor is a small, privileged, and nested software component within Wallet’s TCB. It manages the instantiation and invocation of serverless functions within lightweight protection domains isolated from the guest OS. It performs security-critical tasks such as memory management, function deployment, and attestation.
Within the CVM, Wallet introduces the notion of trusted processes for hosting serverless functions. These processes possess enclave-like features: (i) isolation from the guest OS, (ii) trusted boot, and (iii) remote attestation capabilities. They are categorized into zygotes and trustlets. Zygotes serve as pre-initialized templates containing runtime environments (e.g., Python), while trustlets are lightweight instances, derived from zygotes, to execute functions in an isolated manner. To facilitate inter-function communication, Wallet provides the abstraction of data objects, managed by the trusted monitor ensuring isolation. This design favors reduced boot times, efficient memory use via copy-on-write [5, 53, 116, 52, 96], and optimized data exchange between co-located functions.
@@ 4.4 Design Principles and Primitives
We now present the design principles and primitives of Wallet that address the six key problems, outlined in § 3.
#1: Nested confidential execution. As a countermeasure to the large TCB, the hardware-limited scalability and the costly inter-function communication of CVM-based serverless deployments, Wallet introduces nested confidential execution. Inspired by nested virtualization [24, 83, 32, 109, 164]—not directly applicable to CVMs due to host VMM support requirements [66, 4]—Wallet leverages intra-CVM isolation [1, 72, 110] to partition runtime components [104]. This allows the secure coexistence of untrusted components in lower-privilege compartments, while critical services operate in highly privileged domains, reducing the software TCB.
@@
#2: Decoupled guest OS architecture. Conventional CVMs include the, often bloated [65], guest OS in their TCB. Wallet decouples the guest OS from trusted operations and removes its boot and measurement process from the function invocation path, mirroring ARM TrustZone’s principles, i.e., secure/normal world separation [18, 115], as performed in prior works, such as TLR (Trusted Language Runtime) [133].
After trusted monitor initialization, Wallet boots the untrusted guest OS, totally isolated from function execution environments. This decoupling excludes the guest OS kernel from the TCB, significantly reducing the attack surface and ensuring the guest OS vulnerabilities do not compromise the security of serverless functions. The guest OS handles only auxiliary tasks such as receiving requests and forwarding function invocations to the trusted monitor, maintaining compatibility with existing serverless frameworks. This reduces startup times compared to CVM-based deployments where cold starts involve full guest OS initialization and measurement.
#3: Trusted process templates. Existing confidential serverless deployments must initialize each function instance from scratch. Startup optimizations (e.g., snapshotting) are impractical [65] due to memory confidentiality (e.g., encryption) requirements. To this end, Wallet introduces trusted processes as the core abstraction for executing serverless functions.
Building on this execution model, Wallet implements its differential attestation protocol, which incrementally builds cumulative trust chains by reusing zygote and function measurements, requiring only the measurement of mutable components (e.g., input/output) during invocation, thus reducing attestation latency for both standalone and chained functions.
#4: Dynamically loadable LibOS architecture. To minimize function startup latency and its TCB while supporting diverse workloads (e.g., Python, Node.js), Wallet employs a dynamically loadable LibOS architecture [166, 137], adopting concepts of unikernel-based systems [100, 124, 166]. However, instead of requiring developers to build custom LibOS images, Wallet dynamically loads lightweight LibOS-based runtimes tailored to specific functions at runtime. By providing the language runtimes that directly call a function, common serverless functions can be immediately used inside Wallet.
The LibOS forms the core of Wallet’s trusted processes, providing them with essential runtime functionalities. By dynamically loading minimal runtimes rather than full OSes or containers, Wallet reduces initialization overhead. This design allows on-demand instantiation of lightweight, pre-measured templates and avoids redundant software measurement steps during function invocation, further lowering startup latency without increasing developer effort.
#5: Confidential networking architecture. CVM I/O stacks [117, 44] suffer from performance penalties due to data copies, VM exits, and high CPU utilization for cryptographic operations [92, 112, 62]. Wallet’s key insight is to leverage its function density (§ 3.5) and provide a monitor-mediated, high-performance networking architecture for CVMs.
Wallet leverages function co-location and employs a data-centric I/O approach [161] to optimize function communication via secure data objects, building on insights from prior work (e.g., CAP-VM [135], Nephele [99], Pheromone [161]). Wallet’s high function density increases co-location opportunities, and thus, it can also benefit from data-centric schedulers [2, 82, 161] that effectively co-locate related functions.
In Wallet, data objects serve as secure endpoints for trustlets to exchange data without traversing conventional I/O stacks. The trusted monitor manages data object allocation and page table permissions, granting appropriate access to producers and consumers, and ensures secure data exchange while maintaining isolation between functions. This design eliminates unnecessary cryptographic operations and redundant data copies, reducing I/O overheads. For functions that cannot be co-located, Wallet falls back to standard CVM I/O, ensuring compatibility with existing serverless frameworks.
5 Design
5.1 Trusted Monitor
The trusted monitor is Wallet’s core component. It maintains strict isolation between each serverless function and the guest OS through hardware-enforced CVM partitioning and page table configurations. It operates at the highest privilege level of the CVM partitioning domains (e.g., VMPL-0).
Trusted process management. Wallet’s trusted monitor manages the CVM resources and preserves the state of the guest OS and trusted processes using process descriptors. A process descriptor includes the values for general-purpose and control registers, address space information (page tables), process ID, CVM-partitioning privilege level, object information (§ 5.4), and process state. The trusted monitor is responsible for allocating resources and scheduling the guest OS and the trusted processes. Internally, the descriptor state is linked to one vCPU state of the CVM, and the trusted monitor uses a special instruction to switch the context (e.g., AP creation call in SEV-SNP [14]). By executing the guest OS and the trusted processes at lower CVM partitioning levels, the trusted monitor governs their execution.
Programming interface. As its execution environment base, Wallet adopts a LibOS architecture [23, 121, 31, 54, 91, 27]. This aligns well with the requirements of trusted processes, where the main code runs at the ring3, and offers users maximum flexibility through a common POSIX-compatible interface with a minimal resource footprint. Wallet tailors the LibOS architecture for serverless execution by allowing the dynamic loading of functions without requiring the recompilation of other components. Furthermore, the LibOS provides a data-centric I/O [161] API for inter-function communication (§ 5.2) and relies on function provider policies that include workflow information (e.g., function chaining), measurements of the involved zygotes and functions, and, optionally, remote storage encryption keys for functions requiring external storage access.